OpenStudy (anonymous):

which expression is simplified form of square root of h^11 A. h^3 square root h^2 B. h^10 square root h C. h^5 square root h D. h square root h^5

OpenStudy (hoblos):

sqrt(h^11) = sqrt(h^10 * h) = h^5 sqrt(h)